OpenStudy (anonymous):

what is the distance between point I(-8,2)and point j (4,7)

jigglypuff314 (jigglypuff314):

\[distance formula: \sqrt{(x _{2}-x _{1})^{2} + (y _{2}-y _{1})^{2}}\]when given points (x1, y1) and (x2, y2) so plug in points (-8, 2) and (4 , 7)

jigglypuff314 (jigglypuff314):

so simplify \[\sqrt{(4+ 8)^{2} + (7-2)^{2}}\]

OpenStudy (anonymous):

ok ty so would it be 235 how do I brake that down please

jigglypuff314 (jigglypuff314):

mmm try again? :) 4+8 = 12 7-2 = 5 12^2 = 144 5^2 = 25 144 + 25 = 169 square root of 169 = (answer) :)